OpenStudy (anonymous):

If 2y cos theta=x sin theta and 2x sec theta-y cosec theta=3,then the relation between x and y is

OpenStudy (anonymous):

2y/x=sin theta/cos theta i think

OpenStudy (anonymous):

Let cos theta be u and sin theta be v 2yu = xv --(1) (2x/u) - (y/v) =3 --(2) [because sec theta = (1/cos theta) and cos theta = (1/sin theta) From (2), 2xv - yu = 3uv 2xv - 3uv = yu y = (2xv - 3vu) / u --(3) Sub (3) into (1), 2u[(2xv-3uv)/u] = xv *cancelling u from numerator and denominator* 2(2xv-3uv) = xv 4xv-6uv = xv 3xv = 6uv *cancelling v from both sides* 3x = 6u x = 2u --(4) Sub (4) into (3) y = [2(2u)v - 3vu)] / u y = (4vu -3vu) / u y = vu / u y = v --(5) From (4), x = 2u x = 2cos theta and y = sin theta

OpenStudy (anonymous):

In this case we use the method of substitution to simplify the equations to solve this problem, hope it helps! :)

OpenStudy (anonymous):

nope .i need the equation ................@HatcrewS .something went wrong here.

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@uskidoll

OpenStudy (dumbcow):

@janani1 @HatcrewS is correct, you just need to continue where she left off to get it in terms of only"x" and "y" x = 2 cos ---> x/2 = cos y = sin \[\sin^2 \theta + \cos^2 \theta = 1\] \[\rightarrow y^2 + (\frac{x}{2})^2 = 1\] \[y^2 + \frac{x^2}{4} = 1\] this is the relation between x and y ---> an ellipse with major axis of 4 and minor axis of 2

OpenStudy (lichking):

nice one @HatcrewS